OpenStudy (anonymous):

Leia is baking a cake. The recipe calls for 3/4 cups of flour for the frosting and 5/8 cups of flour for the cake. When you add it you get 8/12 right?

OpenStudy (anonymous):

questions.

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

3/4 + 5/8 = (3*2+5)/8 = 11/8

OpenStudy (anonymous):

gt bro hi

OpenStudy (king):

11/8

OpenStudy (anonymous):

she-hi! where are you from?

OpenStudy (anonymous):

m from india u??

OpenStudy (anonymous):

sheg -hi! where are you from?

OpenStudy (anonymous):

originally from india (hyderabad). now in US. Email me - gtumuluri@yahoo.com.

OpenStudy (anonymous):

yeah sure buddy wt do u do?

OpenStudy (anonymous):

I am a technology executive.

OpenStudy (anonymous):

k great